# Gravitational Wave Signatures from Periodic Orbits around a non-commutative inspired black hole surrounded by quintessence

### v1

@@ I Introduction
The advent of gravitational wave astronomy, marked by the groundbreaking detection of gravitational waves by LIGO and Virgo in 2015, has opened a new frontier in our exploration of the universe LIGOScientific:2016aoc; LIGOScientific:2016vbw; LIGOScientific:2016vlm; LIGOScientific:2016emj. These spacetime ripples, predicted by Einstein’s general theory of relativity, offer a unique observational window into the most energetic and violent cosmic events, such as binary black holes and binary neutron star mergers. Beyond these cataclysmic phenomena, the study of particle trajectories around black holes provides a powerful theoretical framework for probing the intricate dynamics of strong gravitational fields. Among these trajectories, periodic orbits are particularly significant because of their role in addressing fundamental challenges in astrodynamics. The analysis of periodic orbits sheds light not only on the stability of celestial systems and the complex interactions between black holes and their surrounding matter, but also provides fundamental insights into generic orbital dynamics Levin:2008mq; Levin:2009sk; Misra:2010pu; Babar:2017gsg. All generic orbits around black holes can be considered as minor deviations from periodic orbits Levin:2008mq. The study of periodic orbits and their gravitational wave emissions is also of particular interest because of their potential observational applications in future space-based gravitational wave detectors.
@@
Black holes with stellar mass or neutron stars are often found in close orbits around supermassive black holes. Such binary systems are known as the extreme mass ratio inspiral (EMRI), being one of the most critical targets of future space-based gravitational detectors, such as Taiji Hu:2017mde, Tianqin TianQin:2015yph; Gong:2021gvw, LISA Danzmann:1997hm; Schutz:1999xj; Gair:2004iv; LISA:2017pwj; Maselli:2021men, etc. Given that the energy carried away by the orbital motion of the lower mass object is an exceedingly small fraction of the total energy of the system, the time it takes for the smaller mass object to spiral around the supermassive black hole can span several years. During this process, the orbital dynamics of the smaller-mass object can be well approximated by periodic orbits.
@@
To characterize periodic orbits, classifying massive particles is beneficial for understanding the dynamics of black hole mergers Levin:2008mq. The primary concept of this classification scheme is that a dynamic system can be understood by studying its periodic orbits. To be exact, there are three topological integers indexing all closed orbits around a black hole, representing scaling ($z$), rotation ($\omega$), and vertex ($\nu$) behaviors, respectively. Under this taxonomy, extensive research has been carried out on periodic orbits within various black hole spacetimes, to mention a few, including those of Schwarzschild and Kerr Levin:2008ci; Levin:2009sk; Bambhaniya:2020zno; Rana:2019bsn, charged black hole Misra:2010pu, naked singularities Babar:2017gsg, Kerr-Sen black holes Liu:2018vea, and hairy black holes in Horndeski’s theory Lin:2023rmo. For the studies of periodic orbits in other black holes, see refs. Yao:2023ziq; Lin:2022llz; Chan:2025ocy; Wang:2022tfo; Lin:2023eyd; Haroon:2025rzx; Habibina:2022ztd; Zhang:2022psr; Lin:2022wda; Gao:2021arw; Lin:2021noq; Deng:2020yfm; Tu:2023xab; Zhou:2020zys; Gao:2020wjz; Deng:2020hxw; Azreg-Ainou:2020bfl; Wei:2019zdf; Pugliese:2013xfa; Zhang:2022zox; Healy:2009zm; Wang:2025wob; Alloqulov:2025bxh; Wei:2025qlh and references therein. The gravitational wave emissions from the periodic orbits of a large number of black hole spacetimes have also been studied; see refs. Tu:2023xab; Yang:2024lmj; Shabbir:2025kqh; Junior:2024tmi; Zhao:2024exh; Jiang:2024cpe; Yang:2024cnd; Meng:2024cnq; Li:2024tld; QiQi:2024dwc; Haroon:2025rzx; Alloqulov:2025ucf; Wang:2025hla; Lu:2025cxx; Zare:2025aek; Gong:2025mne; Li:2025sfe; Choudhury:2025qsh; Chen:2025aqh; Deng:2025wzz; Li:2025eln; Zahra:2025tdo and references therein.
@@
In this paper, we investigate the gravitational wave emission from the periodic orbital motion of a test particle around a black hole in the context of noncommutative theory. The primary purpose of this article is to study the periodic orbital behaviors of a particle surrounding a black hole in non-commutative geometry and their corresponding gravitational wave radiations. We explore how noncommutative effects affect the behavior of orbits and calculate the corresponding gravitational wave radiation. The article is constructed as follows. In Section II, we present a brief review of non-commutative black hole solutions. Then, in Section III, we discuss the geodesics of a massive test particle around black holes in noncommutative geometry and study the corresponding periodic orbits. In Section IV, we calculate the gravitational wave radiation of periodic orbits around black holes. Conclusion and discussion are presented in Section V.
@@ II Black holes in non-commutative inspired geometry
[…]
@@
We know that in a commutative spacetime, one has a point mass described by the Dirac delta; however, noncommutative geometry naturally smears point sources over a finite length scale $\sqrt{\Theta}$, replacing the Dirac delta. The non-commutative inspired black hole was first constructed using a Gaussian mass profile (Nicolini:2005vd), and alternative smeared densities have since been proposed (see e.g. (Giri:2006rc; Anacleto:2019tdj; Hamil:2024ppj)). In this work, we adopt a regular mass distribution $\rho_{\Theta}(r)$ (see Sec. II for the explicit form) so that the usual mass parameter $M$ in Eq. (2.2) is replaced by a radially dependent smeared mass $M(r,\Theta)$. This procedure regularises the central singularity at scales $r\lesssim\sqrt{\Theta}$ and recovers the Kiselev/non-commutative limits when $\Theta\to 0$ and $p\to 0$, respectively.
@@
where $\theta_{ab}$ is a $4\times 4$ antisymmetric matrix that defines the fundamental discretisation scale of spacetime. This noncommutativity leads to a generalised uncertainty principle (GUP)
@@
indicating that spacetime points cannot be localized with arbitrary precision. This framework refines semiclassical gravity by incorporating noncommutative effects that are expected to appear in a quantum theory of gravity. Nicolini et al. Nicolini:2005vd first realized this idea by constructing a noncommutative geometry–inspired Schwarzschild black hole as an exact solution of Einstein’s equations with a static, spherically symmetric, Gaussian-smeared matter source. equations for the static, spherically symmetric, Gaussian-smeared matter source.
The uncertainty relation indicates that spacetime cannot be sharply defined, leading to an intrinsic fuzziness Hamil (Hamil:2024ppj). The noncommutative parameter $\Theta$, a small positive constant, measures this fuzziness and sets the scale of the minimal length. Different smeared mass distributions have been discussed in the literature (Giri:2006rc; Anacleto:2019tdj; Nicolini:2005gy; Kumar:2017hgs; Ghosh:2017odw; Ghosh:2020cob). In the present work, we follow the form proposed by Anacleto et al. (Anacleto:2019tdj):
@@
In the limiting case $\Theta\to 0$, the spacetime reduces to the Kiselev black hole surrounded by quintessence (Kiselev:2002dx), while for $p\to 0$, it approaches the noncommutative Schwarzschild solution (Nicolini:2005vd). When both $\Theta\to 0$ and $p\to 0$, the standard Schwarzschild geometry is recovered.
@@ III Periodic orbits
The periodic time-like orbits around a black hole enveloped by quintessence and inspired by noncommutative geometry are covered in this section. Understanding the complex structure of bound trajectories in strong gravitational fields requires an examination of periodic orbits  Levin:2008yp. Let us first consider the motion of a test particle in the spacetime of a black hole. The Lagrangian that governs the dynamics of the particle reads
@@
which governs the radial and polar motion of a test particle in the noncommutative geometry. The study of such orbits provides a natural framework to classify zoom–whirl periodic trajectories and to explore their observational signatures in noncommutative black hole spacetimes Levin:2008yp; Levin:2009sk; Misra:2010pu.
@@
This equation is convenient for numerical integration and helps in understanding the stability of circular orbits, as well as how they evolve into periodic or zoom-whirl trajectories in strong gravitational fields Levin:2008yp; Levin:2009sk; Misra:2010pu; Chandrasekhar:1985kt.
@@
We adopt the recipe introduced by Levin and Perez-Giz Levin:2008mq, which classifies all periodic orbits around black holes using a triplet of integers $(z,w,v)$, corresponding to the zoom, whirl, and vertex behavior of the trajectory. In their scheme, a periodic orbit returns to its initial conditions after a finite time, which requires that the ratio of radial to azimuthal frequencies be a rational number. Because a rational one can approximate any irrational number, periodic orbits can effectively represent generic bound trajectories around black holes. The Levin and Perez-Giz Levin:2008mq recipe has been successfully applied to various black holes, including Schwarzschild and Kerr geometries Levin:2009sk; Misra:2010pu; Babar:2017gsg, and provides a useful framework for studying the corresponding gravitational radiation from these orbits.
According to the taxonomy of Levin:2008mq, the ratio $q$ between the two frequencies $\omega_{r}$ and $\omega_{\phi}$ of oscillations in the $r$-motion and $\phi$-motion, respectively, in terms of three integers $(z,w,v)$ as
@@
The integers $(z,w,v)$ each have different geometric meanings. The zoom number $z$ counts the larger circles in an orbit, while the whirl number $w$ counts the small loops near the center. The vertex number $v$ tells us if the particle moves through the orbit’s vertices in a clockwise or counterclockwise direction. To avoid degeneracy, $z$ and $v$ should be relatively prime Levin:2008mq. The parameter $q$ indicates the degree to which the periapsis deviates from that of a simple ellipse, allowing us to understand the orbit’s shape. This framework also considers the order in which the orbital paths or segments are traced. Together, all these numbers help describe the complex behavior of periodic orbits. The ratio $\frac{\omega_{\phi}}{\omega_{r}}$ equals $\Delta\phi/(2\pi)$, where $\Delta\phi=\oint d\phi$ is the total equatorial angle during a period in $r$, and this must be a multiple of the total number of $2\pi$. With the geodesic equations for non-commutative black holes, we can calculate $q$ as follows:
@@
The behavior of $q$ as $E$ and $L_{z}$ vary can be found in Azreg-Ainou:2020bfl. In Figs. 1 and 2, we illustrate the periodic orbits of non-commutative black holes for different combinations of integers $(z,w,v)$. It is worth mentioning that we set $M=1$ for simplicity in the figures. The value of $z$ determines the number of blades in the orbit’s shape. The larger $z$ values correspond to larger blade profiles and increasingly complex trajectories.
@@ IV Gravitational Radiation in non-commutative geometry
In this section, we present a preliminary analysis of the gravitational radiation emitted by a test particle moving in periodic orbits around SMBHs modelled by our noncommutative geometry–inspired solution. The Extreme Mass Ratio Inspirals (EMRIs), consisting of a stellar-mass compact object orbiting an SMBH, are among the most promising sources for future space-based GW detectors such as LISA, Taiji, and TianQin LISA:2017pwj; Gong:2021gvw; Hu:2017mde. The GWs generated by these systems encode detailed information about the strong-field dynamics and the underlying spacetime geometry of the central black hole. A possible observational test of quantum gravity-inspired models could be provided if the smaller body travels on a periodic orbit in a noncommutative spacetime, where spacetime coordinates obey nontrivial commutation relations, and its emitted waveform carries imprints of noncommutative effects.
The analysis of gravitational waveforms from EMRIs is typically carried out using the adiabatic approximation, which assumes that the inspiral timescale is much longer than the orbital period Hughes:1999bq; Barack:2003fp. The motion of the smaller object can be described as a series of geodesics in the SMBH’s background metric, as its energy and angular momentum change slowly in this regime. For short-term orbital evolution Isoyama:2021jjd, the radiation response, or back-reaction of the emitted GWs on the particle’s motion, is ignored at leading order.
@@
We employ a waveform model that offers a practical framework for computing the GWs emitted by periodic orbits in a black hole spacetime, following the approach developed in Babak:2006uv – often referred to as the numerical kludge scheme– that proceeds in two main steps. First, the motion of the small compact object is obtained by numerically integrating the geodesic equations in the background spacetime of the black hole. In the second step, the corresponding gravitational waveform is constructed using the standard quadrupole formula for gravitational radiation. This semi-relativistic approximation has been widely used to model GW signals from EMRIs and provides a powerful tool for analyzing the dynamics of the orbit, the properties of the central black hole, and possible environmental effects Barack:2003fp; Gair:2004iv; Hughes:2000ssa. For a metric perturbation $h_{ij}$ representing the GW and a symmetric, trace-free (STF) mass quadrupole moment $I_{ij}$, the quadrupole formula takes the form
@@
where $A=c^{4}D_{L}/(2G)$, $G=c=1$, and $D_{L}$ is the luminosity distance to the source. By numerically solving the geodesic equations, one obtains the trajectory $Z_{i}(t)$ of the small object in the curved spacetime of the supermassive black hole, which is then used to compute the GW signal. For a particle of mass $m$ moving along a trajectory $Z^{i}(t)$, the quadrupole moment $I_{ij}$ is defined as Thorne:1980ru
@@
The choice of coordinate system plays a key role in both the computation and interpretation of gravitational waveforms. While the geodesic equations are usually solved in the coordinates $(r,\theta,\phi)$, the resulting waveform is conveniently expressed in a detector-adapted Cartesian coordinates $(X,Y,Z)$, which simplifies the analysis of the signal measured by a gravitational-wave detector. The transformation is given by Babak:2006uv
@@
where $v_{i}$ and $a_{i}$ denote the velocity and acceleration components of the small object, respectively, and $A=c^{4}D_{L}/(2G)$ with $G=c=1$. This formalism adheres to the conventional method of numerical kludge waveforms Barack:2003fp; Gair:2004iv; Babak:2006uv; Hughes:2000ssa, which provides an effective and physically consistent approach to approximating EMRI waveforms.
To analyze the gravitational-wave signal as observed by a detector, it is convenient to introduce a detector-adapted Cartesian coordinate system $(X,Y,Z)$, centred on the black hole and oriented with respect to the source frame $(x,y,z)$ by the inclination angle $\iota$ and the longitude of pericentre $\zeta$ Babak:2006uv; Barack:2003fp; Gair:2004iv. This transformation facilitates the projection of the waveform onto the detector frame, enabling the computation of the observable GW polarisations. The unit vectors of the detector frame in the $(x,y,z)$ coordinates are:
@@
where the components are Babak:2006uv
@@
The resulting gravitational waveforms, represented by the two independent components $h_{+}$ and $h_{\times}$, exhibit a characteristic alternating pattern. During the portions of the orbit where the trajectory extends outward in a highly eccentric fashion (the zoom phases), the waveform amplitude remains relatively low. These intervals are followed by short, intense bursts of radiation associated with the nearly circular segments of the trajectory (the whirl phases). The number of low-amplitude intervals corresponds to the number of zoom segments, while the number of intense bursts matches the number of whirls in the orbit. The numerical results obtained from Eqs. (4.10) and (4.4) are shown in Figs. 3 and 4, which clearly display the distinct “zoom” and “whirl” features of the GW signal from periodic orbits in EMRIs, reflecting the orbital dynamics of the small object over one complete cycle Levin:2008mq; Barack:2003fp; Babak:2006uv; Gair:2004iv.
@@
The frequency spectra are obtained by applying a discrete Fourier transform (DFT) to the time-domain gravitational waveforms, converting the signal into the frequency domain. This transformation enables a detailed examination of the signal’s frequency content, revealing how the particle’s periodic orbital motion modulates the structure of the emitted gravitational waves (see Figs. 5 and 6). The dominant frequencies of these signals lie primarily in the millihertz range, making them especially relevant for space-based detectors such as LISA LISA:2017pwj; Barack:2003fp; Gair:2004iv; Babak:2006uv, which are designed to detect low-frequency gravitational waves from EMRIs. The characteristic spectra for different periodic orbits, labelled by the triplet $(z,w,v)$, are displayed in Figs. 5 and 6.

### v2

@@ I Introduction
The advent of gravitational wave astronomy, marked by the groundbreaking detection of gravitational waves by LIGO and Virgo in 2015, has opened a new frontier in our exploration of the universe Abbott and others (2016c, a, d, b). These spacetime ripples, predicted by Einstein’s general theory of relativity, offer a unique observational window into the most energetic and violent cosmic events, such as binary black holes and binary neutron star mergers. Beyond these cataclysmic phenomena, the study of particle trajectories around black holes provides a powerful theoretical framework for probing the intricate dynamics of strong gravitational fields. Among these trajectories, periodic orbits are particularly significant because of their role in addressing fundamental challenges in astrodynamics. The analysis of periodic orbits sheds light not only on the stability of celestial systems and the complex interactions between black holes and their surrounding matter, but also provides fundamental insights into generic orbital dynamics Levin and Perez-Giz (2008); Levin (2009); Misra and Levin (2010); Babar et al. (2017). All generic orbits around black holes can be considered as minor deviations from periodic orbits Levin and Perez-Giz (2008). The study of periodic orbits and their gravitational wave emissions is also of particular interest because of their potential observational applications in future space-based gravitational wave detectors.
@@
Black holes with stellar mass or neutron stars are often found in close orbits around supermassive black holes (SMBHs). Such binary systems are known as the extreme mass ratio inspiral (EMRI), being one of the most critical targets of future space-based gravitational detectors, such as Taiji Hu and Wu (2017), Tianqin Luo and others (2016); Gong et al. (2021), LISA Danzmann (1997); Schutz (1999); Gair et al. (2004); Amaro-Seoane and others (2017); Maselli et al. (2022), etc. Analyzing the signals of the gravitational waveforms allows for precise measurements of the compact object’s orbital motion and the black hole’s gravitational field, offering key insights into the evolution of the Universe and strong-field gravity Bian and others (2026); Ni (2024). Given that the energy carried away by the orbital motion of the lower mass object is an exceedingly small fraction of the total energy of the system, the time it takes for the smaller mass object to spiral around the supermassive black hole can span several years. During this process, the orbital dynamics of the smaller-mass object can be well approximated by periodic orbits.
@@
A systematic classification of periodic orbits for massive particles provides valuable insight into the dynamical processes involved in black hole mergers Levin and Perez-Giz (2008). The primary concept of this classification scheme is that a dynamic system can be understood by studying its periodic orbits. To be exact, there are three topological integers indexing all closed orbits around a black hole, representing scaling ($z$), rotation ($\omega$), and vertex ($\nu$) behaviors, respectively. Under this taxonomy, extensive research has been carried out on periodic orbits within various black hole spacetimes, to mention a few, including those of Schwarzschild and Kerr Levin and Grossman (2009); Levin (2009); Bambhaniya et al. (2021); Rana and Mangalam (2019), charged black hole Misra and Levin (2010), naked singularities Babar et al. (2017), Kerr-Sen black holes Liu et al. (2019), and hairy black holes in Horndeski’s theory Lin and Deng (2023b). For the studies of periodic orbits in other black holes, see refs. Yao and Li (2023); Lin and Deng (2022a); Chan and Lim (2025); Wang et al. (2022); Lin and Deng (2023a); Haroon and Zhu (2025); Habibina and Ramadhan (2023); Zhang and Xie (2022b); Lin and Deng (2022b); Gao and Deng (2021); Lin and Deng (2021); Deng (2020a); Tu et al. (2023); Zhou and Xie (2020); Gao and Deng (2020); Deng (2020b); Azreg-Aïnou et al. (2020); Wei et al. (2019); Pugliese et al. (2017); Zhang and Xie (2022a); Healy et al. (2009); Wang et al. (2025b); Alloqulov et al. (2025a); Wei et al. (2025) and references therein. The gravitational wave emissions from the periodic orbits of a large number of black hole spacetimes have also been studied; see refs. Tu et al. (2023); Yang et al. (2025b); Shabbir et al. (2025); Junior et al. (2025); Zhao et al. (2025); Jiang et al. (2024); Yang et al. (2025a); Meng et al. (2025); Li et al. (2024); Qi et al. (2024); Haroon and Zhu (2025); Alloqulov et al. (2025b); Wang et al. (2025a); Lu and Zhu (2025); Zare et al. (2025); Gong et al. (2025); Li and Kuang (2025); Choudhury et al. (2025); Chen and Yang (2025); Deng et al. (2025); Li et al. (2025); Zahra et al. (2025) and references therein.
@@
In this paper, we investigate the gravitational wave emission from the periodic orbital motion of a test particle around a black hole surrounded by quintessence in the context of noncommutative theory. The primary purpose of this article is to study the periodic orbital behaviors of a particle surrounding a black hole in non-commutative geometry and their corresponding gravitational wave radiations. We explore how noncommutative effects affect the behavior of orbits and calculate the corresponding gravitational wave radiation. The article is constructed as follows. In Section II, we present a brief review of non-commutative black hole solutions. Then, in Section III, we discuss the geodesics of a massive test particle around black holes in noncommutative geometry and study the corresponding periodic orbits. In Section IV, we calculate the gravitational wave radiation of periodic orbits around black holes. Conclusion and discussion are presented in Section V.
@@ II Black holes in non-commutative inspired geometry
where $M$ denotes the black hole mass, $\omega$ is the quintessence equation-of-state parameter (commonly taken in the range $1<\omega<-1/3$, and $p$ is a positive normalization constant characterising the quintessence distribution. The last term in the above Eq. (2.2) follows the Kiselev ansatz for a static, anisotropic fluid that models quintessence; in this construction, the energy density of the quintessence fluid scales with radius and is related to the parameter $p$ (see e.g. (Kiselev, 2003)). In the limit $p\to 0$ the metric (2.1) reduces to the Schwarzschild solution. To set the stage, consider a spacetime in which quintessence, characterized by its density $\rho_{q}$, is present with
@@
We know that in a commutative spacetime, one has a point mass described by the Dirac delta; however, noncommutative geometry naturally smears point sources over a finite length scale $\sqrt{\Theta}$, replacing the Dirac delta. The non-commutative inspired black hole was first constructed using a Gaussian mass profile (Nicolini et al., 2006), and alternative smeared densities have since been proposed (see e.g. (Giri, 2007; Anacleto et al., 2020; Hamil and Lütfüoğlu, 2024)). In this work, we adopt a regular mass distribution $\rho_{\Theta}(r)$ so that the usual mass parameter $M$ in Eq. (2.2) is replaced by a radially dependent smeared mass $M(r,\Theta)$. This procedure regularises the central singularity at scales $r\lesssim\sqrt{\Theta}$ and recovers the Kiselev/non-commutative limits when $\Theta\to 0$ and $p\to 0$, respectively.
@@
where $\theta_{ab}$ is a $4\times 4$ antisymmetric matrix that defines the fundamental discretisation scale of spacetime. This $4\times 4$ antisymmetric matrix has $6$ independent components (similar to an electromagnetic field tensor). To replace it by a single scalar noncommutativity scale, we need a Lorentz-invariant quantity. The natural invariant of an antisymmetric 2-form is defined as $\theta_{ab}\theta^{ab}$, which is represented by $\Theta$ in this paper. This noncommutativity leads to a generalised uncertainty principle (GUP)
@@
indicating that spacetime points cannot be localized with arbitrary precision. This framework refines semiclassical gravity by incorporating noncommutative effects that are expected to appear in a quantum theory of gravity. Nicolini et al. Nicolini et al. (2006) first realized this idea by constructing a noncommutative geometry–inspired Schwarzschild black hole as an exact solution of Einstein’s equations with a static, spherically symmetric, Gaussian-smeared matter source.
The uncertainty relation indicates that spacetime cannot be sharply defined, leading to an intrinsic fuzziness Hamil (Hamil and Lütfüoğlu, 2024). The noncommutative parameter $\Theta$, a small positive constant, measures this fuzziness and sets the scale of the minimal length. Different smeared mass distributions have been discussed in the literature (Giri, 2007; Anacleto et al., 2020; Nicolini, 2005; Kumar and Ghosh, 2017; Ghosh, 2018; Ghosh and Maharaj, 2021; Araújo Filho, 2025; Araújo Filho et al., 2025; Filho et al., 2024). In the present work, we follow the form proposed by Anacleto et al. (Anacleto et al., 2020):
@@
In the limiting case $\Theta\to 0$, the spacetime reduces to the Kiselev black hole surrounded by quintessence (Kiselev, 2003), while for $p\to 0$, it approaches the noncommutative Schwarzschild solution (Nicolini et al., 2006). When both $\Theta\to 0$ and $p\to 0$, the standard Schwarzschild geometry is recovered.
@@ III Periodic orbits
The periodic time-like orbits around a black hole enveloped by quintessence and inspired by noncommutative geometry are covered in this section. Understanding the complex structure of bound trajectories in strong gravitational fields requires an examination of periodic orbits  Levin and Perez-Giz (2009). Let us first consider the motion of a test particle in the spacetime of a black hole. The Lagrangian that governs the dynamics of the particle reads
@@
which governs the radial and polar motion of a test particle in the noncommutative geometry. The study of such orbits provides a natural framework to classify zoom–whirl periodic trajectories and to explore their observational signatures in noncommutative black hole spacetimes Levin and Perez-Giz (2009); Levin (2009); Misra and Levin (2010).
@@
This equation is convenient for numerical integration and helps in understanding the stability of circular orbits, as well as how they evolve into periodic or zoom-whirl trajectories in strong gravitational fields Levin and Perez-Giz (2009); Levin (2009); Misra and Levin (2010); Chandrasekhar (1985).
@@
We adopt the recipe introduced by Levin and Perez-Giz Levin and Perez-Giz (2008), which classifies all periodic orbits around black holes using a triplet of integers $(z,w,v)$, corresponding to the zoom, whirl, and vertex behavior of the trajectory. In their scheme, a periodic orbit returns to its initial conditions after a finite time, which requires that the ratio of radial to azimuthal frequencies be a rational number. Because a rational one can approximate any irrational number, periodic orbits can effectively represent generic bound trajectories around black holes. The Levin and Perez-Giz Levin and Perez-Giz (2008) recipe has been successfully applied to various black holes, including Schwarzschild and Kerr geometries Levin (2009); Misra and Levin (2010); Babar et al. (2017), and provides a useful framework for studying the corresponding gravitational radiation from these orbits.
According to the taxonomy of Levin and Perez-Giz (2008), the ratio $q$ between the two frequencies $\omega_{r}$ and $\omega_{\phi}$ of oscillations in the $r$-motion and $\phi$-motion, respectively, in terms of three integers $(z,w,v)$ as
@@
The integers $(z,w,v)$ each have different geometric meanings. The zoom number $z$ counts the larger circles in an orbit, while the whirl number $w$ counts the small loops near the center. The vertex number $v$ tells us if the particle moves through the orbit’s vertices in a clockwise or counterclockwise direction. To avoid degeneracy, $z$ and $v$ should be relatively prime Levin and Perez-Giz (2008). The parameter $q$ indicates the degree to which the periapsis deviates from that of a simple ellipse, allowing us to understand the orbit’s shape. This framework also considers the order in which the orbital paths or segments are traced. Together, all these numbers help describe the complex behavior of periodic orbits. The ratio $\frac{\omega_{\phi}}{\omega_{r}}$ equals $\Delta\phi/(2\pi)$, where $\Delta\phi=\oint d\phi$ is the total equatorial angle during a period in $r$, and this must be a multiple of the total number of $2\pi$. With the geodesic equations for non-commutative black holes, we can calculate $q$ as follows:
@@
The behavior of $q$ as $E$ and $L_{z}$ vary can be found in Azreg-Aïnou et al. (2020). In Figs. 1 and 2, we illustrate the periodic orbits of non-commutative black holes for different combinations of integers $(z,w,v)$. It is worth mentioning that we set $M=1$ for simplicity in the figures. In addition, we fixed $\omega=-2/3$ and $p=0.001$ for all calculations. The value of $z$ determines the number of blades in the orbit’s shape. The larger $z$ values correspond to larger blade profiles and increasingly complex trajectories.
@@ IV Gravitational Radiation in non-commutative geometry
In this section, we present a preliminary analysis of the gravitational radiation emitted by a test particle moving in periodic orbits around SMBHs modelled by our noncommutative geometry–inspired solution. The Extreme Mass Ratio Inspirals (EMRIs), consisting of a stellar-mass compact object orbiting an SMBH, are among the most promising sources for future space-based GW detectors such as LISA, Taiji, and TianQin Amaro-Seoane and others (2017); Gong et al. (2021); Hu and Wu (2017). The GWs generated by these systems encode detailed information about the strong-field dynamics and the underlying spacetime geometry of the central black hole. A possible observational test of quantum gravity-inspired models could be provided if the smaller body travels on a periodic orbit in a noncommutative spacetime, where spacetime coordinates obey nontrivial commutation relations, and its emitted waveform carries imprints of noncommutative effects.
The analysis of gravitational waveforms from EMRIs is typically carried out using the adiabatic approximation, which assumes that the inspiral timescale is much longer than the orbital period Hughes (2000); Barack and Cutler (2004). The motion of the smaller object can be described as a series of geodesics in the SMBH’s background metric, as its energy and angular momentum change slowly in this regime. For short-term orbital evolution Isoyama et al. (2022), the radiation response, or back-reaction of the emitted GWs on the particle’s motion, is ignored at leading order.
@@
We employ a waveform model that offers a practical framework for computing the GWs emitted by periodic orbits in a black hole spacetime, following the approach developed in Babak et al. (2007) – often referred to as the numerical kludge scheme– that proceeds in two main steps. First, the motion of the small compact object is obtained by numerically integrating the geodesic equations in the background spacetime of the black hole. In the second step, the corresponding gravitational waveform is constructed using the standard quadrupole formula for gravitational radiation. This semi-relativistic approximation has been widely used to model GW signals from EMRIs and provides a powerful tool for analyzing the dynamics of the orbit, the properties of the central black hole, and possible environmental effects Barack and Cutler (2004); Gair et al. (2004); Hughes (2001). For a metric perturbation $h_{ij}$ representing the GW and a symmetric, trace-free (STF) mass quadrupole moment $I_{ij}$, the quadrupole formula takes the form
@@
where $A=c^{4}D_{L}/(2G)$, $G=c=1$, and $D_{L}$ is the luminosity distance to the source. By numerically solving the geodesic equations, one obtains the trajectory $Z_{i}(t)$ of the small object in the curved spacetime of the supermassive black hole, which is then used to compute the GW signal. For a particle of mass $m$ moving along a trajectory $Z^{i}(t)$, the quadrupole moment $I_{ij}$ is defined as Thorne (1980)
@@
The choice of coordinate system plays a key role in both the computation and interpretation of gravitational waveforms. While the geodesic equations are usually solved in the coordinates $(r,\theta,\phi)$, the resulting waveform is conveniently expressed in a detector-adapted Cartesian coordinates $(X,Y,Z)$, which simplifies the analysis of the signal measured by a gravitational-wave detector. The transformation is given by Babak et al. (2007)
@@
where $v_{i}$ and $a_{i}$ denote the velocity and acceleration components of the small object, respectively, and $A=c^{4}D_{L}/(2G)$ with $G=c=1$. This formalism adheres to the conventional method of numerical kludge waveforms Barack and Cutler (2004); Gair et al. (2004); Babak et al. (2007); Hughes (2001), which provides an effective and physically consistent approach to approximating EMRI waveforms.
To analyze the gravitational-wave signal as observed by a detector, it is convenient to introduce a detector-adapted Cartesian coordinate system $(X,Y,Z)$, centred on the black hole and oriented with respect to the source frame $(x,y,z)$ by the inclination angle $\iota$ and the longitude of pericentre $\zeta$ Babak et al. (2007); Barack and Cutler (2004); Gair et al. (2004). This transformation facilitates the projection of the waveform onto the detector frame, enabling the computation of the observable GW polarisations. The unit vectors of the detector frame in the $(x,y,z)$ coordinates are:
@@
where the components are Babak et al. (2007)
@@
The resulting gravitational waveforms, represented by the two independent components $h_{+}$ and $h_{\times}$, exhibit a characteristic alternating pattern. During the portions of the orbit where the trajectory extends outward in a highly eccentric fashion (the zoom phases), the waveform amplitude remains relatively low. These intervals are followed by short, intense bursts of radiation associated with the nearly circular segments of the trajectory (the whirl phases). The number of low-amplitude intervals corresponds to the number of zoom segments, while the number of intense bursts matches the number of whirls in the orbit. The numerical results obtained from Eqs. (4.10) and (4.4) are shown in Figs. 3 and 4, which clearly display the distinct “zoom” and “whirl” features of the GW signal from periodic orbits in EMRIs, reflecting the orbital dynamics of the small object over one complete cycle Levin and Perez-Giz (2008); Barack and Cutler (2004); Babak et al. (2007); Gair et al. (2004).
@@
where, $\tilde{h}(f)=\int h(t)e^{-2\pi ift}dt$ is the one-sided Fourier amplitude (positive frequencies only). This $h_{c}(f)$ is the intrinsic, polarization-combined strain (no detector antenna factors) and satisfies signal to noise ratio (SNR) as $\mathrm{SNR}^{2}=\int(h_{c}^{2}/fS_{n})d\ln f$ Depies (2009); Larson et al. (2000). The frequency spectra are obtained by applying a discrete Fourier transform (DFT) to the time-domain gravitational waveforms, converting the signal into the frequency domain. This transformation enables a detailed examination of the signal’s frequency content, revealing how the particle’s periodic orbital motion modulates the structure of the emitted gravitational waves (see Figs. 5 and 6). The dominant frequencies of these signals lie primarily in the millihertz range, making them especially relevant for space-based detectors such as LISA Amaro-Seoane and others (2017); Barack and Cutler (2004); Gair et al. (2004); Babak et al. (2007), which are designed to detect low-frequency gravitational waves from EMRIs. The characteristic spectra for different periodic orbits, labelled by the triplet $(z,w,v)$, are displayed in Figs. 5 and 6. In the above analysis, Using a frequency resolution of $\approx 0.001/2$ Hz and a maximum frequency of $0.01$ Hz (see Figs. 5 and 6), the total duration of the signal is $T\approx 2000$s. For a sampling frequency of $0.01$ Hz, this corresponds to $N=20$ samples over the interval, yielding a time spacing of $\Delta t\approx 100$s per sample in Figs. 3 and 4.